A well designed and safe workplace can contribute greatly to increas significance level (alpha) of 0.10 ed productivity, The accompanying data on maximun weight of lift MAWL in kg) was randomly selected from ation of healty males age 18-30. Does the followin g data suggest that the population mean is equal to 25 kg? The population standard deviation is 545. Use a 25 23 27 21 24 22 23 o a Tone population b Zone population C. T two popusation o d. Z two population
Answer #1

I HOPE ITS HELP FULL TO YOU HAVE ANY DOUBTS PLS COMMENTS BELLOW...

I WILL BE THERE TO HELP YOU .....PLS RATE THUMBS UP....!!!!

SOL ;;-

AS FOR GIVEN DATA..

The accompanying data on maximun weight of lift (MAWL in kg) was randomly selected from the population of healty males age 18 – 30. Does the following data suggest that the population mean is equal to 25 kg? The population standard deviation is 5.45. Use a significance level (alpha) of 0.10.

25 23 27 21 24 22 23

SOL:

Option ( A ) is correct. We will use Z one population because population standard deviation is known

H0: u = 25

H1: u is not equal to 25

Sample mean = (25+23+27+21+24+22+23)/7 = 23.57

Z = ((23.57-25)/5.45)* Sqrt (7) = -0.6942

Critical value are -1.645 and 1.645

Since test statistic lies between critical values we failed to reject the null hypothesis and conclude that mean is not significantly different

THANK YOU...!!!